OpenStudy (anonymous):

3. The population, P, of a lakeside town with a large number of seasonal residents can be modeled using the function P(t)=5000sin(30(t−7))+8000 , where t is the number of months after New Year’s Day. d) When is the population about 10 000?

OpenStudy (anonymous):

5000sin(30(t−7))+8000=10000 5000 Sin(30(t-7))=2000 Sin(30(t-7))=2/5

OpenStudy (anonymous):

are there two answers to this question